OpenStudy (anonymous):

\[|Z+ \frac{ 9 }{ Z }| = 6\] ,Then find the Greatest value of |Z| ?

OpenStudy (jamesj):

Easiest way to proceed is the square both sides and see what you get from there

OpenStudy (jamesj):

so ... waiting for you to take a step or engage in the conversation

OpenStudy (anonymous):

couldnt you also just say Z + 9/Z = 6, and ignore the absolute value, because you are looking for the largest value anyways then solve for Z as: Z^2 + 9 = 6z Z^2 - 6x + 9 = 0 and go from there

OpenStudy (jamesj):

that gives you some of the solutions. You also need to consider the case z + 9/z = -6

OpenStudy (anonymous):

@JamesJ that would give answer as |Z| = 3.... But the answer shuld be |Z| = 3 + sqrt18

OpenStudy (anonymous):

Note: Z is a Complex Number

OpenStudy (jamesj):

Squaring both sides \[ \left| z + 9/z \right|^2 = 6^2 \] i.e., \[ (z + 9/z)^2 = 36 \] as |x|^2 = x^2 for all real numbers x. Hence \[ z^2 + 18 + 81/z^2 = 36 \] and therefore \[ z^4 - 18z^2 + 81 = 0 \] Now you have a quadratic equation in z^2 which you can solve... \[ (z^2 - 9)^2 = 0 \] \[ z^2 = 9 \] \[ |z| = 3 \] yes. The other answer you just wrote down doesn't satisfy the original equation, does it?

OpenStudy (jamesj):

Or is z here a complex number?

OpenStudy (anonymous):

Z is a Complex Number

OpenStudy (jamesj):

If z is complex, we're going to have use a different approach.

OpenStudy (jamesj):

right. So in that case, what have you tried?

OpenStudy (anonymous):

No Idea..

OpenStudy (jamesj):

Write z = x +iy as usual. Now notice the original equation is equivalent to |z^2 + 9| = 6|z| Now write out those two expressions in terms of x and y and see what you get. It will be easier to square them.

OpenStudy (jamesj):

i.e., write out \[ |z^2 + 9|^2 = 36 |z|^2 \] for example, \[ |z|^2 = x^2 + y^2 \] You take the next step ....

OpenStudy (anonymous):

(x^2 -y^2 + 9)^2 + (2xy)^2 = 36(x^2 + y^2)

OpenStudy (jamesj):

Yes, now see if you can reduce this somehow

OpenStudy (experimentx):